turkkk:

Determine the solution to the system of equations graphed below and explain your reasoning in complete sentences.

UkuleleGirl:

Could you show a screenshot please?

turkkk:

yes give me a second @ukulelegirl

turkkk:

UkuleleGirl:

I recommend you use this This is Desmos Graphing calculator

UkuleleGirl:

Click on the link.

turkkk:

i clicked on the link but im not sure as to what im supposed to write in my answer box.. I assume that I put both equations into desmos but what action should i take afterwards.

surjithayer:

both the graphs cut at (0,1) the solution is the point which lies on both the functions. Hence the solution is x=? y=?

turkkk:

so i would type "since both graphs cut at (0,1) the answer is x=y:? @surjithayer

UkuleleGirl:

They went offline

turkkk:

oh okay thanks for letting me know

UkuleleGirl:

I'm getting someone else.

turkkk:

okay much appreciated

UkuleleGirl:

Wait wait I think I get it now.

UkuleleGirl:

Okay I understand what they were talking about

turkkk:

Okay I'm listening

UkuleleGirl:

So they said look where the point lies on both of the functions

UkuleleGirl:

I'm looking at the graph and it would only make sense for the intersection of 2 lines at one point

turkkk:

the intersection would be (0,2) correct?

J4ke:

correct the solution is (0,2)

turkkk:

Thanks! I appreciate the help.

J4ke:

g(x)=3x+2 notice that g(x)=y our solution is a point (x,y) so lets apply that we have the point (0,2) y=g(x) y=2 and x=0 2=3(0)+2 So 2=2 the point is a solution apply that for the other equation

surjithayer:

correction they both cut at (0,2) so the solution is x=0,y=2

surjithayer:

yes you are correct. it cuts at (0,2)

surjithayer:

g(x)=3x+2 let g(x)=y y=3x+2 ...(1) f(x)=|x-1|+1 let f(x)=y y=|x-1|+1 \[y=\pm (x-1)+1\] if x>1 either y=x-1+1=x x=3x+2 2x=-2 x=-1 rejected as x>1 or x<1 x-1<0 y=-(x-1)+1=-x+2 -x+2=3x+2 4x=0 x=0/4=0 y=3(0)+2=2 Hence solution is (0,2)